# Modeling complex particle suspensions: perspectives on the rigid multiblob method

### v1

@@ III.2.1 Stochastic particle velocities
Eq. (36) shows that the thermal noise is balanced by the viscous dissipation as required by the fluctuation dissipation balance. Thus the stochastic velocity imposed at the blob level (35) generates Brownian body velocities consistent with the Stokes equation. Equivalently, the square root of the body mobility matrix $\bm{N}^{1/2}=\bm{N}\bm{K}^{T}\bm{M}^{-1/2}$ satisfies the fluctuation dissipation balance with $\bm{N}^{1/2}(\bm{N}^{1/2})^{T}=\bm{N}$.
@@ III.2.2 Drift term
$k_{B}T\partial_{\bm{x}}\cdot\bm{N}=k_{B}T\left\langle\frac{\bm{N}\left(\bm{x}+\delta\bm{W}\right)-\bm{N}\left(\bm{x}\right)}{\delta}\cdot\bm{W}\right\rangle+O(\delta),$ (37)
where $\delta$ a small number. This involves applying the mobility matrix $\bm{N}$, and thus solving the mobility problem (34), evaluated at randomly displaced positions/orientations $\delta\bm{W}$ to a random force/torque vector $k_{B}T\bm{W}$. RFD therefore requires at least one additional mobility solve of (34) per time-step compared to the deterministic setting.
@@ III.3.2 Fast hydrodynamics with grid-based methods
where $\mathcal{J}=\mathcal{S}^{\star}$ is adjoint to the spreading operator.
@@
where $\bm{u}_{\mbox{\scriptsize{th}}}=\sqrt{2k_{B}T}\bm{M}^{1/2}_{\text{IBM}}\bm{Z}(t)$, with $\bm{M}^{1/2}_{\text{IBM}}=\mathcal{J}\mathcal{L}^{-1}\mathcal{D}$.
@@ III.5 The multi-physics multiblob method
Figure 5: Illustration and validation of the RMB method for reactive particles. a) Swimming speed for a Janus sphere as of function of the Damkohler number. Results for resolutions with different number of blobs, $N_{b}$, and comparison with the results of Michelin and Lauga [102]. The results with 42 blobs agree well for high Da numbers and shows an error of about $5\%$ for $\text{Da}\ll 1$; for finer resolutions the agreement is good for all Da numbers. Inset shows a the concentration field around a Janus particle (adapted from [102]). b) Velocities along the $x$-axis for two spheres with uniform surface activity as a function of the gap $d$ between the spheres. The motion is symmetric, thus only the velocity magnitude is shown. The results for discretizations with $N_{b}=12,\,42\text{ and }162$ are compared with multipolar expansions [138] and the exact solution computed in bispherical coordinates. Inset: sketch of the configuration. Figure adapted from [38].
@@ IV.0.1 Active particles
For instance, a bacterium can be modeled as a rigid body with a rigid flagellum attached by inextensible links so that the flagellum is free to rotate around its main axis but otherwise moves with the bacterium body. Within this formalism it is possible to apply equal but opposite torques to the bacterium body and flagellum to mimic the effect of the molecular motor acting on the flagellum, so the whole bacterium swims like a free-force swimmer (see Fig. 4). This approach was used with the RMB method to model the swimming dynamics of $100$ bacteria above a no-slip wall, see Fig. 6b [36]. The bacteria were hydrodynamically attracted by the no-slip wall [146] and formed a large cluster, as shown in the snapshot in Fig. 6b, which eventually broke up. Such a large simulation, with $61,800$ unknowns, was possible because the mobility problem was solved in a moderate number of iterations, because the hydrodynamic interactions were computed with a fast method [114, 140], and because it was possible to use large time step sizes as the inextensible links did not introduce any stiffness into the equations of motion.
@@
Phoretic nanorods. Another problem, studied with the RMB method, where the interplay between hydrodynamic interactions and Brownian motion is important is the dynamics of phoretic nanorods. Bimetallic nanorods can self-propel through electrophoresis in aqueous $\mbox{H}_{2}\mbox{O}_{2}$ solutions [147]. These nanorods, similar to the ones shown in Fig. 1g, are formed by two segments of different metals, typically gold and platinum, and decompose the $\mbox{H}_{2}\mbox{O}_{2}$ with asymmetric reactions on each metallic segment [148]. The chemical reactions produce a flux of ions near the nanomotor surface that drags the fluid and by momentum conservation propels the nanorod in the opposite direction. These nanorods have a tendency to swim against flow currents and also against gravity along tilted walls, i.e. they show rheotaxis and gravitaxis [45, 46]. RMB simulations were used to understand these behaviors. In those simulations the nanorods were discretized with a moderated number of blobs ($\sim 100$). Instead of solving the complex electrochemical reactions an active slip velocity was set by hand near the bimetallic interface, see Fig. 6d.
@@ IV.0.2 Complex colloidal suspensions
Rheology of complex shaped colloids. Some colloids are formed by the aggregation of smaller particles. For example, spherical silica particles can fuse to form complex colloids [2], see Fig. 1b. Suspensions formed by such colloids have an interesting rheology, exhibiting high viscosities at low volume fractions and continuous or discontinuous shear thickening [2]. Modeling such systems with traditional methods is too costly due to the complex shapes of the colloids and the role of Brownian motion in their rheological response. However, within the RMB’s framework each silica particle can be modeled as a single blob which reduces the numerical complexity and cost, see Fig. 7a-b. RMB studies of star colloids suspensions have shown that their viscosity can be quite large at low and moderate volume fractions in agreement with experimental results. A hundred fold increase in the viscosity at 20% volume fraction was observed when the colloids have complex shapes, like hooks, that allow them to link to each other and form a network that resists stress [54]. However, in contrast with experimental results that found a shear thickening response [2], RMB studies found shear thinning [37, 54]. Such difference may come from the lack of friction and lubrication forces in the simulations, although both of them can be incorporated into the RMB formulation.
@@
Friction has already been used in the context of Stokesian Dynamics and it can be included as local pairwise force between blobs of different colloids [151]. Since friction forces do not modify the linear systems discussed in Sec. III their implementation is simple, although, it may introduce stiffness into the equations of motion. Including lubrication corrections modifies the mobility problem (32), however, the additional complexity is small. The idea of including lubrication corrections was first introduced in the Stokesian dynamics method [152, 28] and it has been discussed in the context of rigid assemblies of blobs [43, 33].
@@ V Perspectives
The formulation discussed in Sec. II.1 can be generalized in several ways. For example, the no-slip boundary conditions can be extended to partial-slip or Navier boundary conditions [153]. Navier boundary conditions assume that the flow velocity at solid surfaces does not match the surface velocity but instead slips with a velocity proportional to the local velocity gradient. The proportionality factor between the velocity gradient and the slip velocity is the slip length, which represents the distance across the boundary where the linearly extrapolated flow velocity would vanish. At nanoscales, even small slip lengths can significantly affect the dynamics of suspended particles [154]. Boundary integral formulations for solving the Stokes equations with Navier boundary conditions exist, and the RMB can incorporate them into its formalism [155, 154, 156]. These formulations involve the double-layer operator of the Stokes equations, which is not positive definite. Therefore, the approach used in Sec. III.2 to generate Brownian velocities is not directly applicable, and how to efficiently generate Brownian velocities in this setting remains an open problem.
In this manuscript, we have only discussed problems where the suspended particles are rigid objects. However, blobs can also be used to model deformable objects, such as fibers. Assemblies of rigid and elastic objects can thus be built with blobs without any additional complexity. Elastic interactions can introduce stiffness into the equations of motion, which may require specialized integrators. This approach has been used, for example, to study the coordinated motion of cilia on a spherical microorganism [157].
@@
To simulate large numbers of complex particles, it is necessary to use as few blobs per particle as possible while maintaining reasonable accuracy in the hydrodynamic interactions. Despite the regularization error introduced by the regularized Green’s function (25), which scales with the blob radius $a$ and thus as $N_{b}^{-1/2}$ [38], it is possible to enhance numerical accuracy for a given resolution through a grid optimization procedure. Once the number of blobs per particle and the surface mesh are selected, two additional grid parameters remain to be tuned: the blob radius $a$ and the position of the blobs relative to the actual particle surface. Due to their finite size, blobs do not need to lie exactly on the particle surface. The distance of blob centers to the particle surface is described by a geometric parameter $S$, representing the scale of the discretized surface relative to the actual particle surface: when $S=1$, the surface of the grid coincides with the surface of the actual particle; if $S<1$ ($S>1$), the blobs are inside (outside) the surface (see Fig. 8a). In the early works on the multiblob method, this tuning was done empirically and only for the mobility matrix $\bm{N}$ [34, 9]. More recently, Broms et al. [161] formalized the optimization of $\bm{N}$ through a minimization problem. Building on their approach, we proposed a new optimization strategy based on the singular value decomposition of $\widetilde{\bm{N}}$ to match the surface mobility operator as well [38]. Figure 8b shows a comparison of the propulsion speed of a spherical squirmer using the empirically chosen grid [34] and the optimized grid with the exact solution $U_{\mbox{\scriptsize{th}}}$. The optimized grid (triangles) significantly reduces the error compared to the original discretization (disks), enabling the simulation of slip-driven particles with coarse grids while achieving 2–3 digits of accuracy—comparable to that of other methods using much finer grids [156]. This grid optimization strategy is applicable to particles of any shape. One of the key challenges remaining is the development of quadrature rules that can balance, or at least further reduce, the regularization error already present in the continuous setting, with the goal of achieving a convergence rate faster than $\sim N_{b}^{-1/2}$ for the rigid multiblob method.
@@ Acknowledgments
The authors would like to pay tribute to Aleks Donev, who was a key contributor to the rigid multiblob method. B.D. acknowledges support from the French National Research Agency (ANR), under award ANR-20-CE30-0006. F.B.U. acknowledges funding by the Basque Government through the BERC 2022-2025 program and by the Ministry of Science, Innovation and Universities: BCAM Severo Ochoa accreditation CEX2021-001142-S/MICIN/AEI/10.13039/501100011033.

### v2

@@ III.2.1 Stochastic particle velocities
Eq. (40) shows that the thermal noise is balanced by the viscous dissipation as required by the fluctuation-dissipation theorem (11). Thus the stochastic velocity imposed at the blob level (39) generates Brownian body velocities consistent with the Stokes equation. Equivalently, the square root of the body mobility matrix $\bm{N}^{1/2}=\bm{N}\bm{K}^{T}\bm{M}^{-1/2}$ satisfies the fluctuation-dissipation theorem with $\bm{N}^{1/2}(\bm{N}^{1/2})^{T}=\bm{N}$.
@@ III.2.2 Drift term
$k_{B}T\partial_{\bm{x}}\cdot\bm{N}=k_{B}T\left\langle\frac{\bm{N}\left(\bm{x}+\delta\bm{W}/2\right)-\bm{N}\left(\bm{x}-\delta\bm{W}/2\right)}{\delta}\cdot\bm{W}\right\rangle+O(\delta^{2}),$ (41)
where $\delta$ a small number. This involves applying the mobility matrix $\bm{N}$, and thus solving two mobility problems (38), evaluated at randomly displaced positions/orientations ±$\pm\delta\bm{W}/2$ to a random force/torque vector $k_{B}T\bm{W}$. Since the parameter $\delta$ is small the solution of one mobility problem can be used as initial guess to the second mobility problem reducing the computational cost [41].
@@ III.3.2 Fast hydrodynamics with grid-based methods
where $\mathcal{J}=\mathcal{S}^{T}$ is adjoint to the spreading operator.
@@
The IBM mobility differs from the RPY mobility defined in Eq. (28). However, one can perform a second order Taylor expansion of the flow field inside the kernels to show that $\bm{M}_{\text{IBM}}$ obeys (29) to order $\mathcal{O}\left(a^{4}\right)$. Thus, different kernels define different regularizations that are all essentially equivalent to the RPY mobility [92, 149].
@@
where $\bm{u}_{\mbox{\scriptsize{th}}}=\sqrt{2k_{B}T}\bm{M}^{1/2}_{\text{IBM}}\bm{Z}(t)$, with $\bm{M}^{1/2}_{\text{IBM}}=\mathcal{J}\mathcal{L}^{-1}\mathcal{D}$. To fulfill (49), it is crucial that the averaging and spreading operators are adjoint ($\mathcal{J}=\mathcal{S}^{T}$)  [^3] It is also necessary that the discretization of the Stokes equations preserves some of the original symmetries, specifically $\mathcal{L}^{-1}\mathcal{D}\mathcal{D}^{T}\mathcal{L}^{-1}=\mathcal{L}^{-1}$, which is true in the continuum setting but it is not respected by all discretizations [92]., which is not the case for the method of Regularized Stokeslets [35].
@@ III.5 The multi-physics multiblob method
[…]
@@ IV.0.1 Active particles
For instance, a bacterium can be modeled as a rigid body with a rigid flagellum attached by inextensible links so that the flagellum is free to rotate around its main axis but otherwise moves with the bacterium body. Within this formalism it is possible to apply equal but opposite torques to the bacterium body and flagellum to mimic the effect of the molecular motor acting on the flagellum, so the whole bacterium swims like a free-force swimmer (see Fig. 4). This approach was used with the RMB method to model the swimming dynamics of 100100100 bacteria above a no-slip wall, see Fig. 6b [44]. The bacteria were hydrodynamically attracted by the no-slip wall [163] and formed a large cluster, as shown in the snapshot in Fig. 6b, which eventually broke up. Such a large simulation, with 61,80061,80061 , 800 unknowns, was possible because the mobility problem was solved in a moderate number of iterations, because the hydrodynamic interactions were computed with a fast method [131, 158], and because it was possible to use large time step sizes as the inextensible links did not introduce any stiffness into the equations of motion.
@@
Phoretic nanorods. Another problem, studied with the RMB method, where the interplay between hydrodynamic interactions and Brownian motion is important is the dynamics of phoretic nanorods. Bimetallic nanorods can self-propel through electrophoresis in aqueous $\mbox{H}_{2}\mbox{O}_{2}$ solutions [164]. These nanorods, similar to the ones shown in Fig. 1g, are formed by two segments of different metals, typically gold and platinum, and decompose the $\mbox{H}_{2}\mbox{O}_{2}$ with asymmetric reactions on each metallic segment [165]. The chemical reactions produce a flux of ions near the nanomotor surface that drags the fluid and by momentum conservation propels the nanorod in the opposite direction. These nanorods have a tendency to swim against flow currents and also against gravity along tilted walls, i.e. they show rheotaxis and gravitaxis [57, 58]. RMB simulations were used to understand these behaviors. In those simulations the nanorods were discretized with a moderated number of blobs (100$\sim$ 100∼ 100). Instead of solving the complex electrochemical reactions an active slip velocity was set by hand near the bimetallic interface, see Fig. 6d.
@@ IV.0.2 Complex colloidal suspensions
Rheology of complex shaped colloids. Some colloids are formed by the aggregation of smaller particles. For example, spherical silica particles can fuse to form complex colloids [2], see Fig. 1b. Suspensions formed by such colloids have an interesting rheology, exhibiting high viscosities at low volume fractions and continuous or discontinuous shear thickening [2]. Modeling such systems with traditional methods is too costly due to the complex shapes of the colloids and the role of Brownian motion in their rheological response. However, within the RMB’s framework each silica particle can be modeled as a single blob which reduces the numerical complexity and cost, see Fig. 7a-b. RMB studies of star colloids suspensions have shown that their viscosity can be quite large at low and moderate volume fractions in agreement with experimental results. A hundred fold increase in the viscosity at 20% volume fraction was observed when the colloids have complex shapes, like hooks, that allow them to link to each other and form a network that resists stress [67]. However, in contrast with experimental results that found a shear thickening response [2], RMB studies found shear thinning [45, 67]. Such difference may come from the lack of friction and lubrication forces in the simulations, although both of them can be incorporated into the RMB formulation [41].
@@
Friction has already been used in the context of Stokesian Dynamics and it can be included as local pairwise force between blobs of different colloids [168]. Since friction forces do not modify the linear systems discussed in Sec. LABEL:sec:RMB their implementation is simple, although, it may introduce stiffness into the equations of motion. Including lubrication corrections modifies the mobility problem (36), however, the additional complexity is small. The idea of including lubrication corrections was first introduced in the Stokesian dynamics method [169, 13] and it has been discussed in the context of rigid assemblies of blobs [55, 41].
@@ V Perspectives
The formulation discussed in Sec. II.1 can be generalized in several ways. For example, the no-slip boundary conditions can be extended to partial-slip or Navier boundary conditions [170]. Navier boundary conditions assume that the flow velocity at solid surfaces does not match the surface velocity but instead slips with a velocity proportional to the local velocity gradient. The proportionality factor between the velocity gradient and the slip velocity is the slip length, which represents the distance across the boundary where the linearly extrapolated flow velocity would vanish. At nanoscales, even small slip lengths can significantly affect the dynamics of suspended particles [171]. Boundary integral formulations for solving the Stokes equations with Navier boundary conditions exist [172, 171, 79], and the RMB can incorporate them into its formalism [173]. These formulations involve the double-layer operator of the Stokes equations, which is not positive definite. Therefore, the approach used in Sec. III.2 to generate Brownian velocities is not directly applicable, and how to efficiently generate Brownian velocities in this setting remains an open problem.
In this manuscript, we have only discussed problems where the suspended particles are rigid objects. However, blobs can also be used to model deformable objects, such as fibers. Assemblies of rigid and elastic objects can thus be built with blobs without any additional complexity. Elastic interactions can introduce stiffness into the equations of motion, which may require specialized integrators [174]. This approach has been used, for example, to study the coordinated motion of cilia on a spherical microorganism [175].
@@
To simulate large numbers of complex particles, it is necessary to use as few blobs per particle as possible while maintaining reasonable accuracy in the hydrodynamic interactions. Despite the regularization error introduced by the regularized Green’s function (28), which scales with the blob radius $a$ and thus as $N_{b}^{-1/2}$ [46], it is possible to enhance numerical accuracy for a given resolution through a grid optimization procedure. Once the number of blobs per particle and the surface mesh are selected, two additional grid parameters remain to be tuned: the blob radius $a$ and the position of the blobs relative to the actual particle surface. Due to their finite size, blobs do not need to lie exactly on the particle surface. The distance of blob centers to the particle surface is described by a geometric parameter $S$, representing the scale of the discretized surface relative to the actual particle surface: when $S=1$, the surface of the grid coincides with the surface of the actual particle; if $S<1$ ($S>1$), the blobs are inside (outside) the surface (see Fig. 8a). In the early works on the multiblob method, this tuning was done empirically and only for the mobility matrix $\bm{N}$ [42, 12]. More recently, Broms et al. [127] formalized the optimization of $\bm{N}$ through a minimization problem. Building on their approach, we proposed a new optimization strategy based on the singular value decomposition of $\widetilde{\bm{N}}$ to match the surface mobility operator as well [46]. Figure 8b shows a comparison of the propulsion speed of a spherical squirmer using the empirically chosen grid [42] and the optimized grid with the exact solution $U_{\mbox{\scriptsize{th}}}$. The optimized grid (triangles) significantly reduces the error compared to the original discretization (disks), enabling the simulation of slip-driven particles with coarse grids while achieving 2–3 digits of accuracy—comparable to that of other methods using much finer grids [79]. This grid optimization strategy is applicable to particles of any shape. One of the key challenges remaining is the development of quadrature rules that can balance, or at least further reduce, the regularization error already present in the continuous setting, with the goal of achieving a convergence rate faster than $\sim N_{b}^{-1/2}$ for the rigid multiblob method. An interesting alternative lies in recent developments of the method of fundamental solutions, which, although still much more computationally expensive than the RMB, show promising results in terms of accuracy and scalability [178].
@@ Acknowledgments
The authors would like to pay tribute to Aleks Donev, who was a key contributor to the rigid multiblob method. We also thank the anonymous reviewers for their insightful comments. B.D. acknowledges support from the French National Research Agency (ANR), under award ANR-20-CE30-0006. F.B.U. acknowledges funding by the Basque Government through the BERC 2022-2025 program and by the Ministry of Science, Innovation and Universities: BCAM Severo Ochoa accreditation CEX2021-001142-S/MICIN/AEI/10.13039/501100011033.